OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a1 = 16 and a15= -40? a.-152 b.-148 c.-144 d.-140

OpenStudy (anonymous):

a15=-40 a15=a1+(15-1)d -40=16+(14*d) -56=14*d d=-4 a41=a1+(41-1)d =16+(40)*(-4) =16-160 =-144